OpenStudy (anonymous):

integral of x* sqrt(16 - x^2)

OpenStudy (wolfe8):

This can be written as \[x \left( 16-x ^{2} \right)^{1/2}\]You can use the product rule for this.

OpenStudy (anonymous):

how is that ?

OpenStudy (anonymous):

uv* int(v)du ??

OpenStudy (wolfe8):

This might help http://www.mathsrevision.net/product-and-quotient-rules

OpenStudy (anonymous):

let u = 16 - x^2; du = -2x dx \[-\frac{ 1 }{ 2 } \int\limits_{}^{}u^{1/2} du\]